OpenStudy (anonymous):

I need help with the following question, I am confused! A block of material has a volume of 2 m3. It has a mass of 1 × 10^4 kg. Calculate the density of the material sample and express your results in g/cm3 with appropriate scientific notation.

OpenStudy (anonymous):

Do you know what density is?

OpenStudy (anonymous):

No I don't!

OpenStudy (anonymous):

That is all that I was given.

OpenStudy (anonymous):

Density, as the name suggests, is how dense a thing is. Imagine a basket of oranges. There are 5 oranges in this basket. The basket is 30cm^3 large. The density of the bakery would be 5 oranges per 30cm^3.

OpenStudy (anonymous):

So density is the amount of things in a specified volume. Mathematically, density mass over volume.

OpenStudy (anonymous):

Ok, so how would I calculate the density with a mass of 1 × 10^4 kg, and a volume of 2 m3?

OpenStudy (anonymous):

I don't get how the 2 m3 could be divided by 1 x 10^4 kg

OpenStudy (anonymous):

Actually, its the other way around. Divide the mass by the volume

OpenStudy (anonymous):

So the density is 5000000 g/cm^3?

OpenStudy (anonymous):

Nearly there. Did you convert metres cubed into cm cubed?

OpenStudy (anonymous):

Actually, its the other way around. Divide the mass by the volume

OpenStudy (anonymous):

Yes I did

OpenStudy (anonymous):

10^4 kg = 10000kg = 10000000 g 2m^3 = 2000000cm^3

OpenStudy (anonymous):

Did you do those conversions correctly?

OpenStudy (anonymous):

I believe I did, I got 5000000 g/cm^3. Could you confirm if this is true?

OpenStudy (anonymous):

10000000/2000000=5

OpenStudy (anonymous):

So 5g/cm^3 is the correct answer? I got 5000000g/cm^3. Do you know what I could have done wrong?

OpenStudy (anonymous):

How did you get 5000000g/cm3 in the first place?

OpenStudy (anonymous):

Could you show me your work so I can see what I did wrong?

OpenStudy (anonymous):

Oh I think I know

OpenStudy (anonymous):

Does that look good now?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Thank you for being patient and helping!

OpenStudy (anonymous):

No problem. :)